OpenStudy (anonymous):

passing through (4,-2) and perpendicular to x=5/4y-2. Write in slope intercept form. For my answer i have y=-4/5x+6/5 is this correct?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

x=5/4y-2 x+2=5/4y 5/4y = x+2 y = 4/5x + 2(4/5) y = 4/5x + 8/5

jimthompson5910 (jim_thompson5910):

the slope of y = 4/5x + 8/5 is 4/5 so the slope of x=5/4y-2 is also 4/5

jimthompson5910 (jim_thompson5910):

the perpendicular slope would be -5/4 because you flip the fraction and the sign of 4/5

jimthompson5910 (jim_thompson5910):

so your answer will have a slope of -5/4 not -4/5

OpenStudy (anonymous):

the original slope was 5/4 not 4/5

jimthompson5910 (jim_thompson5910):

you first need to solve for y to get the slope

jimthompson5910 (jim_thompson5910):

5/4 is not the original slope

OpenStudy (anonymous):

soo it would then be -4/5 for perp wouldn't it?

jimthompson5910 (jim_thompson5910):

-5/4 for the perp

jimthompson5910 (jim_thompson5910):

4/5 is the original slope

OpenStudy (anonymous):

okay